OpenStudy (anonymous):

(-3/4) X + 2 = (1/2)X + 6 -3.2 5/4 16 4

OpenStudy (anonymous):

(-3/4) X + 2 = (1/2)X + 6 (put all x's on one side and take all other terms to the other side of the equation) (-3/4) X - (1/2)X = 6 - 2 (-5/4) x = 4 x= -16/5

OpenStudy (anonymous):

thank yew dear :3

OpenStudy (anonymous):

np :)